OpenStudy (anonymous):

Reduce the equation to one of the standard forms, classify the surface. (No answers without explanations/ work please. I want to understand why)

OpenStudy (anonymous):

Sorry, you need an equation: 4y^2 + z^2 - x -16y -4z +20 = 0.

OpenStudy (anonymous):

let's complete the squares -x+ 4(y^2 -4y +4) +(z^2-4z+4)=-20 add 16 and 4 to other side as well -x+ 4(y-2)^2 +(z-2)^2=-20+16+4 4(y-2)^2 +(z-2)^2=x In general ----------- a(y^2)+ b(z^2)=1 tends to represent ellipse but our equation is not equal to 1 but x let's to tracing when z=0 we get 4(y-2)^2 +(0-2)^2=x 4(y-2)^2=x-4 that's parabola, along the z axis it looks like parabola  but on y-z plane, setting x=0 4(y-2)^2 +(0-2)^2=0 this is ellipse, so it would be elliptical- paraboloid